### 14 Mar 2024 v2

# VBART: The Turkish LLM

Abstract
We present VBART, the first Turkish sequence-to-sequence Large Language Models (LLMs) pre-trained on a large corpus from scratch. VBART are compact LLMs based on good ideas leveraged from BART and mBART models and come in two sizes, Large and XLarge. Fine-tuned VBART models surpass the prior state-of-the-art results in abstractive text summarization, title generation, text paraphrasing, question answering and question generation tasks. They allow fine-tuning for future text generation tasks and datasets, carving a new path for Turkish Natural Language Processing (NLP) research. Our work shows that having a pre-trained LLM for Turkish outperforms up to 3x multilingual models, improving existing results and providing efficient models for training and inference. Moreover, we show that our monolingual tokenizer is up to 11x more efficient than multilingual tokenizers. Last but not least, we introduce a method to enlarge an existing pre-trained LLM and question the relevancy of Chinchilla Scaling Law to sequence-to-sequence masked language models. Our fine-tuned models, tokenizer and cleaned vngrs-web-corpus of 135 GB are publicly available at huggingface.co/vngrs-ai.
Keywords:  Large Language Models, LLM, Turkish, BART, sequence-to-sequence, Text Summarization, Title Generation, Text Paraphrasing, Question Generation, Question Answering.
[…]
VBART: The Turkish LLM
Meliksah Turker, Mehmet Erdi Ari, Aydin Han
VNGRS-AI
YTÜ Teknopark B2 103 Davutpaşa, İstanbul, Turkey
{meliksah.turker, erdi.ari, aydin.han}@vngrs.com
Abstract content
1.   Introduction
Natural Language Processing (NLP) research landscape has changed drastically over the last decade. Invention of word embedding methods like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) has been the foundation of transfer learning. They have been succeeded by FastText Bojanowski et al. (2017), which leverages character n-grams, and ELMo Peters et al. (2018), which uses the context of the words for the first time.
Using pre-trained word embeddings for text classification tasks often involved Recurrent Neural Networks (RNNs) such as Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) and Gated Recurrent Unit (GRU) Cho et al. (2014). Democratization of these models has been possible thanks to Deep Learning frameworks like Keras Chollet et al. (2015), Tensorflow Abadi et al. (2015) and PyTorch Paszke et al. (2019).
Tokenization methods such as WordPiece Wu et al. (2016), Byte-Pair Encoding (BPE) Sennrich et al. (2015) and Unigram model Kudo (2018) solved the out-of-vocabulary problem while keeping the vocabulary size feasible.
Transformer Vaswani et al. (2017) architecture and the introduction of BERT Devlin et al. (2018) have made a significant contribution to the field of transfer learning. BERT allowed for achieving state-of-the-art results efficiently.
Following the success of BERT, an Encoder-only model, BART Lewis et al. (2019) and T5 Raffel et al. (2019) models have shown that it is possible to apply the principle of unsupervised pre-training of Masked Language Models to seq2seq tasks for conditional text generation.
They are followed by multilingual versions mBART Liu et al. (2020) and mT5 Xue et al. (2020), pre-training of which are conducted on multiple languages together. This allowed them to be fine-tuned for low-resource languages without suffering the high cost of pre-training a new model for the language from scratch. State-of-the-art results in conditional text generation tasks are obtained Baykara and Güngör (2022b); Safaya et al. (2022); Akyon et al. (2022) for Turkish by fine-tuning them.
Figure 1: Pre-training loss of VBART models. Note that there are two steep drops in loss towards the end of the training. This is due to the reduction in Dropout from 0.10 to 0.05 and then to 0.
Lastly, very large language models like BLOOM Scao et al. (2022), PaLM Chowdhery et al. (2022), GPT-4 OpenAI (2023) have made a significant impact on both the industry and the academy. They achieved state-of-the-art results in English and high-resource languages and decent results for low-resource languages. However, as in the case of ChatGPT OpenAI (2022), distilling very large models results in poor performance Lai et al. (2023); Jiao et al. (2023) for low-resource languages. On the other hand, using very large models is computationally expensive.
It is clear that the performance of multilingual LLMs in low-resource languages is a matter of trade-off between the cost of computation and evaluation metrics. This is because they are not optimal for a specific language since a significant portion of training time and network capacity are spent on other languages. This is especially true for low-resource languages such as Turkish. Hence, a dedicated LLM is needed to obtain state-of-the-art results in an efficient and computationally cheap manner for a given language.
Thus, we present VBART, the first dedicated seq2seq LLM pre-trained for Turkish on a large scale. First, we train VBART-Large from scratch on a cleaned corpus of 135.7 GBs of Turkish text based on the pre-training task of BART and the model architecture of mBART. Then, we create VBART-XLarge by enlarging the prior model by doubling the number of encoder and decoder layers, using the weights of VBART-Large. Both models surpass the previous state-of-the-art in abstractive text summarization, paraphrasing, title generation, question answering and question generation tasks. Our contributions to Turkish NLP research are as follows:
  1. 1.
    SentencePiece Unigram Tokenizer
  2. 2.
    vngrs-web-corpus that consist of 135 GBs of cleaned text
  3. 3.
    VBART LLM Models
  4. 4.
    New state-of-the-art results for text generation tasks in Turkish.
2.   Related Work
The existence of BERTurk Dbmdz (2023), a dedicated pre-trained Encoder only Language Model (LM) for Turkish allowed obtaining state-of-the-art results on text Toprak Kesgin et al. (2023); Ozdemir and Yeniterzi (2020); Köksal and Özgür (2021); Şahinuç et al. (2021) and token Aras et al. (2021); Çarık and Yeniterzi (2022) classification tasks.
In the case of text generation tasks, however, researchers are compelled to either use BERTurk-based hybrid solutions or fine-tune pre-trained multilingual seq2seq LLM like mBART and mT5 since there is no dedicated seq2seq LLM for Turkish.
R1/R2/RL Parameters MLSum TRNews XLSum
 Safaya et al., 2022
Base 120M 40.23/27.23/35.08
mBART50 610M 43.75/30.60/38.47
mT5-Base 580M 44.13/31.09/38.76
 Baykara and Güngör, 2022b
BERT2BERT-32K 248M 41.48/27.23/37.66 41.06/25.60/37.69
mBART25 610M 40.47/26.17/36.22 40.52/25.22/36.80
mT5-Base 580M 42.26/27.81/37.96 41.13/25.75/37.60
Our work
VBART-Large 387M 45.75/32.71/39.86 41.97/28.26/36.69 34.15/17.94/28.03
VBART-XLarge 740M 46.13/33.01/40.42 42.55/28.69/37.42 35.13/18.80/29.18
Table 1: Text Summarization
Text Summarization: Baykara and Güngör (2022a) has built a seq2seq model by combining a pre-trained BERT Encoder and a randomly initialized Transformer Decoder. Baykara and Güngör (2022b) and Safaya et al. (2022) have fine-tuned mBART and mT5 on text summarization datasets.
Title Generation While working on the internet news domain for text summarization, Baykara and Güngör (2022a, b) have created a dataset of news titles and fine-tuned mBART and mT5 models to generate title from the given the news summary.
Question Generation & Answering: Akyon et al. (2022) has fine-tuned mT5 for question generation and answering tasks.
Text Paraphrasing: Bağcı and Amasyali (2021) used pre-trained BERT weights to initialize both encoder and decoder weights to create a seq2seq transformer network for text paraphrasing task.
Even though using BERT weights to initialize encoder weights is a reasonable approach, initializing the decoder network weights randomly or setting the same BERT encoder weights for the decoder is not optimal since it results in a model whose encoder and decoder were not pre-trained together to generate text. Fine-tuning a multilingual seq2seq network is a better approach since the model is pre-trained for text generation as a whole. Yet it is not optimal either, as multilingual models are exposed to a trace amount of Turkish during pre-training, and consequently, most of the training time and learning capacity of the network are used for high-resource languages.
Therefore, both of the approaches so far rely on sub-optimal pre-trained models and underachieve the fine-tuning potential.
3.   Model
3.1.   Tokenizer
Before training a model from scratch, a tokenizer is needed first. We use SentencePiece Kudo and Richardson (2018) Unigram Model Tokenizer that is part of VNLP Turker et al. (2024) library. It was trained on 10 GB of text that consists of random subsets of OSCAR Abadji et al. (2022), OPUS Zhang et al. (2020) and Wikipedia dump corpora. Tokenizer training required 500+ GBs of memory and took 3 hours on 96 CPU cores.
3.2.   Network Architecture
The network is based on mBART architecture rather than BART. Following the finding of mBART authors, this is a conscious design choice and serves to stabilize the FP16 mixed precision training Micikevicius et al. (2017) thanks to post-encoder and post-decoder LayerNorm layers. The only difference from the mBART architecture is that we use sinusoidal positional embeddings. The network follows the original BART-Large configuration with 12 encoder and decoder layers, 16 heads and 1024 model dimension. Having a vocabulary size of 32,000 in this configuration results in 387.6M trainable parameters.
3.3.   Pre-training Task
The pre-training task is sentence permutation combined with span masking where 30% of tokens are masked with span length defined by the Poisson distribution ($\lambda=3.5$), following the BART-Large objective.
3.4.   Training Corpus
Training corpus is made of Turkish sections of OSCAR-2201 Abadji et al. (2022) and mC4 Xue et al. (2020), which contain 10.8M and 87.7 million pages, respectively. They are concatenated to obtain the final training corpus of 98.5M pages in total. However, web-crawled data is often noisy and full of keywords, titles and other non-sentence texts. In order to obtain a higher-quality dataset, these pages are cleaned using a chain of rules and heuristics. See the Appendix for the details of the data cleaning process. The cleaned corpus holds 135.7 GB of space on disk, contains 50.3M pages and is made of 25.33B subword tokens in total. We call this curated dataset vngrs-web-corpus.
3.5.   Data Generator
A data generator is written to feed the model training. It works dynamically so that even when the same text input is read from the corpus, stochastic text noising processes result in a different permutation of sentences and tokens masked. Moreover, in the case of samples that are longer than the context length, a random continuous span of sentences that will fit is selected. This has two benefits. First, no sentence is cut in half. Second, the model gets to see other sections of longer texts as well.
Both encoder and decoder context lengths are set to 1024 with right padding. However, noising of the encoder inputs results in shorter sequences since spans of tokens are replaced by a single mask token. This results in a waste of computation on the encoder side. Hence, the encoder context length is set to 800 during pre-training for efficiency.
From News Summary From News Content
R1/R2/RL Parameters MLSum TRNews XLSum MLSum TRNews XLSum
Baykara and Güngör, 2022b
BERT2BERT-32K 248M 39.35/21.14/37.55 41.87/24.37/40.88
mBART25 610M 34.85/18.03/33.46 37.72/20.99/36.74
mT5-Base 580M 40.77/22.42/38.97 41.87/24.49/40.87
Our work
VBART-Large 387M 45.17/30.49/42.92 43.05/29.20/41.87 42.72/26.79/40.70 39.82/25.88/37.83 39.26/25.45/38.11 37.40/21.09/34.71
VBART-XLarge 740M 45.11/30.57/42.98 43.52/29.56/42.16 42.89/26.92/40.70 39.79/25.88/37.79 39.41/25.61/38.25 37.39/21.20/34.87
Table 2: Title Generation
3.6.   Training
Training is carried out on 8X Nvidia A100-80 GB on AWS for 2.7M steps with a batch size of 256 and took 30 days. Adam Kingma and Ba (2014) optimizer ($\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=1e-6$) along with the learning rate scheduler (20,000 warmup steps) from the original Transformer paper is used. Dropout is set to 0.1 for the first 2.33M steps, reduced to 0.05 for the next 165K steps and finally 0 for the last 205K steps.
Considering that the model context length is 1024 tokens and training is carried out with a batch size of 256 on a dataset that is made of 25.33B tokens, 2.7M steps correspond to 28 epochs. Corresponding training loss can be seen in Figure 1.
3.7.   Model Enlargement
After the pre-training of VBART-Large, we decided to create an XLarge version by doubling the number of encoder and decoder layers while keeping the other configuration intact. Having a smaller budget, VBART-XLarge is initialized by using VBART-Large weights to speed up the training. All weights are interchangeable except for the newer Encoder and Decoder layers. Thus, we set every odd-numbered encoder and decoder layer weights from the Large model while initializing even-numbered layer weights from scratch.
Then VBART-XLarge is pre-trained on the same hardware and the pre-training task for 8 days. Due to time constraints and the increased model size, the number of warmup steps and batch size are reduced to 5,000 and 128, respectively. Thanks to weight transfer, the XLarge model reached the loss values that the Large model had reached in 1 and 20 days in 1.5 hours and 3 days, respectively. In total, the XLarge model is trained for 640K steps, with the first 480K steps with a Dropout value of 0.1, the next 80K steps with a Dropout value of 0.05, and the final 80K steps with a Dropout value of 0.
3.8.   Implementation
SentencePiece Kudo and Richardson (2018) library is used to train the Unigram Model tokenizer. The network, data generator and training are implemented using Tensorflow Abadi et al. (2015) framework. Moreover, trained networks and the tokenizer are converted to corresponding Huggingface implementations in order to use Huggingface Transformers’ Wolf et al. (2019)text generation utilities. They can be accessed on the Huggingface hub  [^1] https://huggingface.co/vngrs-ai along with the vngrs-web-corpus.
Parameters Open Subtitles Tatoeba
Alkurdi et al.,2022 BERTScore-{cased/uncased} / BLEU / Rouge-L / METEOR / TER
mT5-Base Open Subtitles 580M 88.89/91.94/36.40/73.87/72.16/37.58 91.61/93.93/34.74/86.60/84.85/18.23
mT5-Base Tatoeba 88.95/92.08/38.13/68.39/65.87/45.13 91.97/94.20/37.02/84.05/81.59/22.76
Our work
VBART-Large 387M 89.25/92.22/47.78/75.35/73.93/35.74 95.79/96.87/69.86/88.14/86.56/16.93
VBART-XLarge 740M 89.30/92.25/47.29/75.38/74.04/35.72 95.66/96.77/68.21/87.56/86.14/17.51
Table 3: Text Paraphrasing
4.   Experiments
After pre-training, VBART models are fine-tuned on various tasks using Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=1e-6$) with different learning rates varying by the downstream task, applying a linear decay of 0.95. Each task’s evaluation is conducted using the referenced work’s evaluation script for consistency.
4.1.   Text Summarization
Large and XLarge models are fine-tuned on Turkish sections of MLSum Scialom et al. (2020), TRNews Baykara and Güngör (2022b), XLSum Hasan et al. (2021) and Wikilingua Ladhak et al. (2020) datasets for 30 and 20 epochs respectively, with a learning rate of $51e-$ 5. Then they are evaluated on the corresponding test splits. Note that Wikilingua does not contain a test split for Turkish, so it is excluded from the evaluation.
Rouge-1, Rouge-2 and Rouge-L scores are computed, and a comparison with the previous work is reported in Table 1. It is the higher the better for Rouge metrics. It is observed that our models surpass the previous state-of-the-art in these datasets on 5 out of 6 metrics, despite the Large model having significantly fewer parameters. The only exception to this is BERT2BERT-32K model, which achieved better Rouge-L on the TRNews dataset. Moreover, we report the results on the XLSum dataset for the first time in the literature in this work.
4.2.   Title Generation
VBART models are fine-tuned to generate titles from given news and summaries. The cartesian product of model sizes and training data results in 4 distinct models. Large model is fine-tuned for 15 and 25 epochs on news summary and content as input, while XLarge model is fine-tuned for 10 and 15 epochs.
Similar to text summarization, Rouge-1, Rouge-2 and Rouge-L are used for evaluation. Comparison with the earlier work is reported in Table 2. Our models surpass the previous work by far in all metrics and datasets. Since Baykara and Güngör (2022b) did not fine-tune any models to generate titles from the news content, there is no result to compare.
Figure 2: Relative number of tokens spent to encode Turkish text in average. The numbers below show the vocabulary size of each tokenizer. Despite their large vocabulary size, multilingual tokenizers spend up to 96% more tokens than VBART Tokenizer.
4.3.   Text Paraphrasing
VBART models are fine-tuned for 20 and 25 epochs, respectively, on a mixture of OpenSubtitles Lison and Tiedemann (2016), TED Cettolo et al. (2012) and Tatoeba Artetxe and Schwenk (2019) datasets mentioned in the paper published by Alkurdi et al. (2022). The original paper suggests a method to improve the dataset quality. Since the filtered data is not published, our models are trained and evaluated on the unfiltered data. Then, we compare our models to the models that were fine-tuned on unfiltered data. Results are evaluated using BERTScore, BLEU, Rouge-L, Meteor and Translation Error Rate (TER) metrics. BERTScore is computed using cased 222dbmdz/bert-base-turkish-cased last reached in 19 Oct 2023 and uncased 333dbmdz/bert-base-turkish-uncased last reached in 19 Oct 2023 versions. Results are reported in Table 3. Except for TER, which is the lower, the better; for all metrics, the higher, the better.
We compare our models with the two models from the referenced work. Our models outperform both mT5-Base models in every metric, especially by a large margin for BERTScore.
Question Generation Task Question Answering Task
BLEU-1 / BLEU-2 / Rouge-L F1 / EM
Parameters TQuADv1 TQuADv2 XQuAD TQuADv1 TQuADv2 XQuAD
Çağatay Akyön et al., 2022
mT5-Small 300M 37.3/30.1/44.3 39.6/32.9/46.5 21.1/13.8/28.4 63.8/48.5 67.1/50.5 48.8/32.9
mT5-Base 580M 48.4/41.7/53.6 47.6/41.2/53.9 27.9/20.9/35.8 72.1/55.8 71.5/56.2 61.1/43.3
mT5-Large 1.2B 49.8/43.2/55.2 49.1/42.7/54.3 29.3/21.9/37.5 74.7/59.6 73.3/58.4 65.0/46.7
Our work
VBART-Large 387M 49.2/42.4/54.8 50.8/43.8/56.5 30.0/22.8/35.2 73.9/58.5 75.6/59.6 60.2/44.0
VBART-XLarge 740M 49.1/42.2/54.9 51.5/44.7/57.1 30.5/23.1/35.7 75.1/59.4 77.1/61.0 62.4/46.7
Table 4: Question Generation and Answering
4.4.   Question Generation and Answering
Our models are fine-tuned on "Turkish NLP Q&A Dataset" Peker (2018) (will be abbreviated as TQuAD) for three tasks - answer extraction, question generation and question answering- using the methodology described in Çağatay Akyön et al. (2022). TQuAD has two variations, TQuADv1 and TQuADv2, with a large intersection of samples. Moreover, there are some samples that exist in both the train and the evaluation sets. These intersections are reported in the Appendix 5 We convert the prompts to Turkish and process the raw data using the script from the referenced work’s github repository 444 OBSS/turkish-question-generation last reached in 19 Sep 2023..
We fine-tune our Large and XLarge models for 50 and 55 epochs, respectively, with a learning rate of $65e-$ 6. The train set is obtained by concatenating TQuaDv1 and TQuaDv2 and dropping the duplicate samples.
Then we evaluate on the test splits of TQuADv1, TQuADv2 and XQuAD Artetxe et al. (2019) and report the results on Table4. BLEU-1, BLEU2 and Rouge-L are computed for question generation task. F1 and Exact Match (EM) are computed for question answering task. In order to be comparable, we do not drop the intersected test samples from the training dataset, following the original work. Also, note that following Çağatay Akyön et al. (2022), we report metrics up to a single decimal in order to be comparable. Therefore, these results are useful for comparison only and are not realistic enough for the tasks themselves.
Our models outperform mT5-Base and mT5-Small and are comparable to mT5-Large, which is a 1.2B parameters model. Out of 15 metrics, VBART-Large and VBART-XLarge models surpass mT5-Large in 7 and 8 metrics, respectively. Overall, mT5-Large is better on TQuADv1, while ours are better on TquADv2.
5.   Discussion
5.1.   Tokenizer
Although one often thinks of the number of model parameters when comparing computational efficiency, it is equally crucial to be able to represent text using fewer tokens. Tokenizers of multilingual models are trained on multilingual corpora. This has two consequences. First, their vocabulary size is chosen to be large to be comprehensive enough. Second, despite their large vocabulary size, it takes more tokens to encode a text in a language compared to a tokenizer that was trained in the specific language alone.
We compare VBART Tokenizer to mBART, mT5 and OpenAI’s cl100k_base tokenizers. In doing so, we randomly sample 1M pages from vngrs-web-corpus and tokenize each page. Then we count the total number of tokens generated by each tokenizer. Figure 2 shows the relative number of tokens required to encode Turkish text, normalized by VBART Tokenizer. Despite having almost 8x vocabulary size, mBART and mT5 tokenizers spend 19.70% and 40.79% more tokens, respectively. OpenAI’s cl100k_base tokenizer, which is used along with Chat-GPT and GPT-4 models, spends 95.99% more tokens despite its more than 3x vocabulary size.
Moreover, taking vocabulary sizes into account, we can claim that VBART Tokenizer’s representation power is $\frac{119.70}{100}\times\frac{250,054}{32,000}=9.35$ times mBART’s. This is 11.00 and 6.14 for mT5 and OpenAI tokenizers, respectively.
5.2.   Experiments
Fine-tuning results of 4 downstream tasks show that having a dedicated pre-trained model for a language outperforms multilingual mT5 and mBART models, even when the number of parameters is significantly less for the prior. VBART-Large model is able to surpass mBART25, mBART50 and mT5-Base models despite having 37% and 33% less parameters, respectively. Moreover, it is observed in Table 4 that VBART-Large is at par with mT5-Large, which has more than 3x parameters.
5.3.   VBART-Large vs. VBART-XLarge
It is evident that the XLarge model’s improvement over the Large model is small. This is because the XLarge model is pre-trained for far fewer steps than the Large one. The Large model is exposed to 708B tokens, while this is only 84B for the XLarge version, 88% less. Despite that, the XLarge version is able to improve the results of the Large, although with small margins. We hypothesize that the XLarge version can improve significantly when pre-trained for more steps.
5.4.   Chinchilla Scaling Law
Chinchilla scaling law Hoffmann et al. (2022) states that, in the case of single epoch training, an LLM is optimal if there are 20 tokens in the training set per network parameter. Then VBART-Large and VBART-XLarge are $\frac{708B\text{ tokens}}{387.6M\times 20}=91.33$ and $\frac{84B\text{ tokens}\times 28\text{ epochs}}{387.6M\times 20}=10.83$ fold optimal, respectively. Accordingly, Large and XLarge models should have been pre-trained for 30K and 59K steps, respectively. However, we observe significant improvements over the course of pre-training in Figure 1.
Moreover, unlike auto-regressive GPT Radford et al. (2018) models, pre-training objective and the dynamic data generator used in this work provide different input-output pairs every time a page is sampled, augmenting the dataset.
Considering these, it is a matter of question whether the Chinchilla Scaling Law is applicable to encoder-decoder models or models with dynamic data generators with a pre-training objective other than next token prediction.
5.5.   Future Work
This work can be extended in multiple ways for the monolingual setting. The most trivial direction is further model enlargement by increasing 24 layers of the XLarge model to 36 layers and hitting 1B parameters. Others are pre-training another Turkish model based on a different architecture, using a different pre-training objective such as T5, or on a larger dataset.
Beyond language-specific directions, there are two research directions that can be explored. First, the model enlargement technique proposed in this work can be exclusively studied. This is a promising direction that can considerable amount of training costs. Second, the applicability of Chinchilla Scaling Law to any other than decoder-only models, the pre-training task of which is next-token prediction, is an open question. Despite the recent popularity of decoder-only models, it would be beneficial to have a heuristic to determine the pre-training configuration of an encoder-decoder model.
6.   Conclusion
In this work, we presented the first sequence-to-sequence LLMs for the Turkish language. All of the components; dataset, tokenizer and models are prepared and trained for Turkish from scratch. We showed our tokenizer’s ability of compact representation over multilingual ones, compared the two models proposed in this work, introduced a method to enlarge an existing LLM and questioned the relevancy of Chinchilla Scaling Law to encoder-decoder models. Achieving new state-of-the-art results for abstractive text summarization, paraphrasing, title generation, question answering and question generation, we showed that monolingual pre-trained models surpass multilingual ones when fine-tuned for downstream tasks, despite having significantly fewer parameters.
Acknowledgements
We thank Amazon Web Services for funding the pre-training of the models.
7.   Bibliographical References
[…]
  • Abadi et al. (2015) Martín Abadi, Ashish Agarwal, Paul Barham, Eugene Brevdo, Zhifeng Chen, Craig Citro, Greg S. Corrado, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Ian Goodfellow, Andrew Harp, Geoffrey Irving, Michael Isard, Yangqing Jia, Rafal Jozefowicz, Lukasz Kaiser, Manjunath Kudlur, Josh Levenberg, Dandelion Mané, Rajat Monga, Sherry Moore, Derek Murray, Chris Olah, Mike Schuster, Jonathon Shlens, Benoit Steiner, Ilya Sutskever, Kunal Talwar, Paul Tucker, Vincent Vanhoucke, Vijay Vasudevan, Fernanda Viégas, Oriol Vinyals, Pete Warden, Martin Wattenberg, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2015. TensorFlow: Large-scale machine learning on heterogeneous systems. Software available from tensorflow.org.
  • Abadji et al. (2022) Julien Abadji, Pedro Ortiz Suarez, Laurent Romary, and Benoît Sagot. 2022. Towards a Cleaner Document-Oriented Multilingual Crawled Corpus. arXiv e-prints, page arXiv:2201.06642.
  • Akyon et al. (2022) Fatih Cagatay Akyon, Ali Devrim Ekin Cavusoglu, Cemil Cengiz, Sinan Onur Altinuc, and Alptekin Temizel. 2022. Automated question generation and question answering from Turkish texts. Turkish Journal of Electrical Engineering and Computer Sciences.
  • Alkurdi et al. (2022) Besher Alkurdi, Hasan Yunus Sarioglu, and Mehmet Fatih Amasyali. 2022. Semantic similarity based filtering for Turkish paraphrase dataset creation. In Proceedings of the 5th International Conference on Natural Language and Speech Processing (ICNLSP 2022), pages 119–127, Trento, Italy. Association for Computational Linguistics.
  • Aras et al. (2021) Gizem Aras, Didem Makaroğlu, Seniz Demir, and Altan Cakir. 2021. An evaluation of recent neural sequence tagging models in turkish named entity recognition. Expert Systems with Applications, 182:115049.
  • Artetxe et al. (2019) Mikel Artetxe, Sebastian Ruder, and Dani Yogatama. 2019. On the cross-lingual transferability of monolingual representations. CoRR, abs/1910.11856.
  • Artetxe and Schwenk (2019) Mikel Artetxe and Holger Schwenk. 2019. Massively multilingual sentence embeddings for zero-shot cross-lingual transfer and beyond. Transactions of the Association for Computational Linguistics, 7:597–610.
  • Baykara and Güngör (2022a) Batuhan Baykara and Tunga Güngör. 2022a. Abstractive text summarization and new large-scale datasets for agglutinative languages turkish and hungarian. Language Resources and Evaluation, 56(3):973–1007.
  • Baykara and Güngör (2022b) Batuhan Baykara and Tunga Güngör. 2022b. Turkish abstractive text summarization using pretrained sequence-to-sequence models. Natural Language Engineering, pages 1–30.
  • Bağcı and Amasyali (2021) Ahmet Bağcı and Mehmet Fatih Amasyali. 2021. Comparison of turkish paraphrase generation models. In 2021 International Conference on INnovations in Intelligent SysTems and Applications (INISTA), pages 1–6.
  • Bojanowski et al. (2017) Piotr Bojanowski, Edouard Grave, Armand Joulin, and Tomas Mikolov. 2017. Enriching word vectors with subword information. Transactions of the association for computational linguistics, 5:135–146.
  • Çarık and Yeniterzi (2022) Buse Çarık and Reyyan Yeniterzi. 2022. A twitter corpus for named entity recognition in turkish. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 4546–4551.
  • Cettolo et al. (2012) Mauro Cettolo, Christian Girardi, and Marcello Federico. 2012. Wit3: Web inventory of transcribed and translated talks. In Proceedings of the Conference of European Association for Machine Translation (EAMT), pages 261–268.
  • Cho et al. (2014) Kyunghyun Cho, Bart Van Merriënboer, Caglar Gulcehre, Dzmitry Bahdanau, Fethi Bougares, Holger Schwenk, and Yoshua Bengio. 2014. Learning phrase representations using rnn encoder-decoder for statistical machine translation. arXiv preprint arXiv:1406.1078.
  • Chollet et al. (2015) François Chollet et al. 2015. Keras. https://keras.io.
  • Chowdhery et al. (2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. 2022. Palm: Scaling language modeling with pathways. arXiv preprint arXiv:2204.02311.
  • Dbmdz (2023) Dbmdz. 2023. dbmdz/bert-base-turkish-uncased. Online; accessed 07-Aug-2023.
  • Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv preprint arXiv:1810.04805.
  • Hasan et al. (2021) Tahmid Hasan, Abhik Bhattacharjee, Md. Saiful Islam, Kazi Mubasshir, Yuan-Fang Li, Yong-Bin Kang, M. Sohel Rahman, and Rifat Shahriyar. 2021. XL-sum: Large-scale multilingual abstractive summarization for 44 languages. In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021, pages 4693–4703, Online. Association for Computational Linguistics.
  • Hochreiter and Schmidhuber (1997) Sepp Hochreiter and Jürgen Schmidhuber. 1997. Long short-term memory. Neural computation, 9(8):1735–1780.
  • Hoffmann et al. (2022) Jordan Hoffmann, Sebastian Borgeaud, Arthur Mensch, Elena Buchatskaya, Trevor Cai, Eliza Rutherford, Diego de Las Casas, Lisa Anne Hendricks, Johannes Welbl, Aidan Clark, et al. 2022. Training compute-optimal large language models. arXiv preprint arXiv:2203.15556.
  • Jiao et al. (2023) Wenxiang Jiao, Wenxuan Wang, JT Huang, Xing Wang, and ZP Tu. 2023. Is chatgpt a good translator? yes with gpt-4 as the engine. arXiv preprint arXiv:2301.08745.
  • Kingma and Ba (2014) Diederik P Kingma and Jimmy Ba. 2014. Adam: A method for stochastic optimization. arXiv preprint arXiv:1412.6980.
  • Köksal and Özgür (2021) Abdullatif Köksal and Arzucan Özgür. 2021. Twitter dataset and evaluation of transformers for turkish sentiment analysis. In 2021 29th Signal Processing and Communications Applications Conference (SIU), pages 1–4. IEEE.
  • Kudo (2018) Taku Kudo. 2018. Subword regularization: Improving neural network translation models with multiple subword candidates. arXiv preprint arXiv:1804.10959.
  • Kudo and Richardson (2018) Taku Kudo and John Richardson. 2018. Sentencepiece: A simple and language independent subword tokenizer and detokenizer for neural text processing. arXiv preprint arXiv:1808.06226.
  • Ladhak et al. (2020) Faisal Ladhak, Esin Durmus, Claire Cardie, and Kathleen McKeown. 2020. Wikilingua: A new benchmark dataset for cross-lingual abstractive summarization. arXiv preprint arXiv:2010.03093.
  • Lai et al. (2023) Viet Dac Lai, Nghia Trung Ngo, Amir Pouran Ben Veyseh, Hieu Man, Franck Dernoncourt, Trung Bui, and Thien Huu Nguyen. 2023. Chatgpt beyond english: Towards a comprehensive evaluation of large language models in multilingual learning. arXiv preprint arXiv:2304.05613.
  • Lewis et al. (2019) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Ves Stoyanov, and Luke Zettlemoyer. 2019. Bart: Denoising sequence-to-sequence pre-training for natural language generation, translation, and comprehension. arXiv preprint arXiv:1910.13461.
  • Lison and Tiedemann (2016) Pierre Lison and Jörg Tiedemann. 2016. Opensubtitles2016: Extracting large parallel corpora from movie and tv subtitles.
  • Liu et al. (2008) Fei Tony Liu, Kai Ming Ting, and Zhi-Hua Zhou. 2008. Isolation forest. In 2008 eighth ieee international conference on data mining, pages 413–422. IEEE.
  • Liu et al. (2020) Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020. Multilingual denoising pre-training for neural machine translation. Transactions of the Association for Computational Linguistics, 8:726–742.
  • Micikevicius et al. (2017) Paulius Micikevicius, Sharan Narang, Jonah Alben, Gregory Diamos, Erich Elsen, David Garcia, Boris Ginsburg, Michael Houston, Oleksii Kuchaiev, Ganesh Venkatesh, et al. 2017. Mixed precision training. arXiv preprint arXiv:1710.03740.
  • Mikolov et al. (2013) Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013. Efficient estimation of word representations in vector space. arXiv preprint arXiv:1301.3781.
  • OpenAI (2022) OpenAI. 2022. Chatgpt.
  • OpenAI (2023) OpenAI. 2023. Gpt-4 technical report. ArXiv, abs/2303.08774.
  • Ozdemir and Yeniterzi (2020) Anil Ozdemir and Reyyan Yeniterzi. 2020. Su-nlp at semeval-2020 task 12: Offensive language identification in turkish tweets. In Proceedings of the Fourteenth Workshop on Semantic Evaluation, pages 2171–2176.
  • Paszke et al. (2019) Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Kopf, Edward Yang, Zachary DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai, and Soumith Chintala. 2019. Pytorch: An imperative style, high-performance deep learning library. In Advances in Neural Information Processing Systems 32, pages 8024–8035. Curran Associates, Inc.
  • Peker (2018) Peker. 2018. Turkish nlp q&a dataset.
  • Pennington et al. (2014) Jeffrey Pennington, Richard Socher, and Christopher D Manning. 2014. Glove: Global vectors for word representation. In Proceedings of the 2014 conference on empirical methods in natural language processing (EMNLP), pages 1532–1543.
  • Peters et al. (2018) Matthew E. Peters, Mark Neumann, Mohit Iyyer, Matt Gardner, Christopher Clark, Kenton Lee, and Luke Zettlemoyer. 2018. Deep contextualized word representations.
  • Radford et al. (2018) Alec Radford, Karthik Narasimhan, Tim Salimans, Ilya Sutskever, et al. 2018. Improving language understanding by generative pre-training. OpenAI.
  • Raffel et al. (2019) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. 2019. Exploring the limits of transfer learning with a unified text-to-text transformer. arXiv e-prints.
  • Safaya et al. (2022) Ali Safaya, Emirhan Kurtuluş, Arda Goktogan, and Deniz Yuret. 2022. Mukayese: Turkish nlp strikes back. In Findings of the Association for Computational Linguistics: ACL 2022, pages 846–863.
  • Şahinuç et al. (2021) Furkan Şahinuç, Çağri Toraman, and Aykut Koç. 2021. Topic detection based on deep learning language model in turkish microblogs. In 2021 29th signal processing and communications applications conference (SIU), pages 1–4. IEEE.
  • Scao et al. (2022) Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, Matthias Gallé, et al. 2022. Bloom: A 176b-parameter open-access multilingual language model. arXiv preprint arXiv:2211.05100.
  • Scialom et al. (2020) Thomas Scialom, Paul-Alexis Dray, Sylvain Lamprier, Benjamin Piwowarski, and Jacopo Staiano. 2020. Mlsum: The multilingual summarization corpus. arXiv preprint arXiv:2004.14900.
  • Sennrich et al. (2015) Rico Sennrich, Barry Haddow, and Alexandra Birch. 2015. Neural machine translation of rare words with subword units. arXiv preprint arXiv:1508.07909.
  • Toprak Kesgin et al. (2023) Himmet Toprak Kesgin, Muzaffer Kaan Yuce, and Mehmet Fatih Amasyali. 2023. Developing and evaluating tiny to medium-sized turkish bert models. arXiv e-prints, pages arXiv–2307.
  • Turker et al. (2024) Meliksah Turker, Erdi Ari, and Aydin Han. 2024. Vnlp: Turkish nlp package. arXiv preprint arXiv:2403.01309.
  • Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. Advances in neural information processing systems, 30.
  • Wolf et al. (2019) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, et al. 2019. Huggingface’s transformers: State-of-the-art natural language processing. arXiv preprint arXiv:1910.03771.
  • Wu et al. (2016) Yonghui Wu, Mike Schuster, Zhifeng Chen, Quoc V Le, Mohammad Norouzi, Wolfgang Macherey, Maxim Krikun, Yuan Cao, Qin Gao, Klaus Macherey, et al. 2016. Google’s neural machine translation system: Bridging the gap between human and machine translation. arXiv preprint arXiv:1609.08144.
  • Xue et al. (2020) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2020. mt5: A massively multilingual pre-trained text-to-text transformer. arXiv preprint arXiv:2010.11934.
  • Zhang et al. (2020) Biao Zhang, Philip Williams, Ivan Titov, and Rico Sennrich. 2020. Improving massively multilingual neural machine translation and zero-shot translation.
  • Çağatay Akyön et al. (2022) Fatih Çağatay Akyön, Ali Devrim Ekin Çavuşoğlu, Cemil Cengiz, Sinan Onur Altınuç, and Alptekin Temizel. 2022. Automated question generation and question answering from turkish texts. Turkish Journal of Electrical Engineering and Computer Sciences, 30(5):1931–1940.
Appendix A Data Cleaning
Data cleaning is conducted on two levels: page cleaning and sentence cleaning. T5 Raffel et al. (2019) authors describe their data cleaning process in detail. In this work, they are leveraged and extended for a cleaner dataset.
A.1.   Page Cleaning
A.1.1.   Anomaly Detection
Some of the pages are filled with Search Engine Optimization (SEO) targeted, repeating keywords, titles and other entities that do not make a sequence of sentences that belong to a context. In order to get rid of these pages in an unsupervised way, Isolation Forest anomaly detection algorithm Liu et al. (2008) is used on five heuristics that separate the mentioned bad pages from the good ones. They are:
  • Mean sentence length on the page
  • Standard deviation of sentence length on the page
  • Maximum sentence length on the page
  • Ratio of short sentences(less than 4 words) to the number of sentences in the page
  • Ratio of uppercase characters to the number of characters on the page
Then, the model is trained and inferred on the whole dataset with default configuration to obtain the anomaly scores.
A.1.2.   Rule-based Cleaning
After obtaining anomaly scores, final page filtering is applied by removing pages
  • whose anomaly score is below 0.05
  • that contains bad/naughty words
  • that contains "lorem ipsum"
  • whose language probability is lower than 0.85 (applicable for OSCAR corpus only)
This results in a 19.81% reduction in the number of pages.
A.2.   Sentence Cleaning
Among the remaining pages, each page is split into sentences, and each sentence is discarded if any of the following is True:
  • It is an empty string
  • It does not end with sentence termination punctuation
  • It contains curly bracket "{" or "}"
  • It contains "JavaScript"
  • It contains "gizlilik ve çerezler" (privacy and cookies)
  • It contains "|"
  • The number of words in the sentence is less than 4 or more than 50
  • The longest word in the sentence is longer than 30 characters, which is the 0.995 quantile
  • Capital letters account for more than half of the sentence
  • 1/3 of the characters are numeric
  • It does not contain any punctuation
  • It contains too many duplicate words; that is, the average number of duplicate words is greater than 2
A.3.   Finalization of the Dataset
As the last step, pages with less than 5 sentences are cleaned, which further reduces the number of pages by 28.69M. In the end, 50.3M pages consisting of 25.33B subword tokens remain to form the 135.7GB dataset.
Appendix B Question Generation & Answering Dataset Intersections
(a) Raw (JSON)
xquad-eval
tquad2-eval
tquad2-train
tquad1-eval
tquad1-train
0 194 7641 194
tquad1-eval
0 882 194
tquad2-train
86 194
tquad2-eval
0
(b) Processed (Script output)
xquad-eval
tquad2-eval
tquad2-train
tquad1-eval
tquad1-train
0 520 20329 520
tquad1-eval
0 2343 520
tquad2-train
227 520
tquad2-eval
0
Table 5: Data intersections
Working on question-answering and question-generation tasks, we have observed that there are intersected samples in the splits that exist in other splits as well, including between train and evaluation splits.
In order to understand the extent and cause of the intersections, we have examined raw and processed versions, conducting a cross-examination. We report the results in Table 4(a) and Table 4(b). Cell values are the number of samples intersected between the corresponding datasets. The numbers under the dataset names indicate the total number of samples in the dataset.